OpenStudy (anonymous):

Solve the equation. 2(3f – 2) = 50 how do i do this??

OpenStudy (espex):

\[2(3f – 2) = 50 \rightarrow 3f-2 = 50/2 \rightarrow 3f = 25 + 2 \rightarrow f = 27/3 \rightarrow f=9\]